# DESI-Like Hubble Expansion From Staged Symmetry Breaking

### v1

@@ IV.2 SIDM Cross Section
We include the closed-form expression for the matrix element in the Appendix. We employ the tree-level cross section instead of partial wave analysis or the simpler Born approximation because the tree-level cross section captures physics that the other two do not. While partial wave analysis has been used to capture non-perturbative effects for a Yukawa interaction [63], we have vectors interacting via a scalar, rather than fermions. Massive vector fields differ from fermions because unlike fermions, they are bosons with three distinct polarizations, which affects the cross section. The Born approximation offers a simpler method for computing a cross section from a Yukawa potential, though it neglects the $u$-channel diagram, as well as the interference between $t$ and $u$-channels [64], which our calculation incorporates. (See Figure 2 for a depiction of the $t$ and $u$-channel processes.) One can view neglecting the $u$-channel as a simplifying approximation, though it dramatically changes $d\sigma/d\Omega$. A process that includes only the $t$-channel yields a differential cross section that peaks at $\theta=0$ but not $\theta=\pi$, whereas the true differential cross section peaks at both $\theta=0$ and $\theta=\pi$, producing a U-shaped plot. This can in turn yield a noticeable impact on the momentum transfer cross section [64], so it is preferable to use the full tree-level cross section, when possible, even though this is more complicated to compute.
@@ V Conclusion
We find that the space of models that can produce similar behavior to that desired by Chen and Loeb for explaining the DESI results is heavily constrained by basic consistency and renormalizability requirements. Many seemingly convenient models are undesirable as quantum field theories. Designing a natural, renormalizable model that produces the desired behavior is somewhat difficult, though at least one model has potential promise. This model also naturally yields a possible SIDM candidate as a consequence of gauge invariance. Future work could fit the model parameters to DESI data, incorporate three-body decays, study effects on structure formation, consider production mechanisms for the dark matter, and more thoroughly investigate the SIDM candidate. While there are a few possibilities for production, one mechanism could involve inflaton decay [65], or a coupling of a dark vector field to the inflaton field [66].
We emphasize that even if future work finds our fiducial model to be disfavored, this study is still valuable for mapping out the space of renormalizable, natural models that have the potential to realize the behavior desired by Chen and Loeb. This is important for finding a pathway towards explaining the DESI trends without incorporating unphysical constructions.
As an aside, it is important to note that to achieve $w=1/2$, which is frequently used by [18], one would need an operator of the form $\lambda\phi^{6}$. This is non-renormalizable, so such a term is problematic. One could try to achieve an effective $\phi^{6}$ from $\chi\phi^{3}$ after integrating out a heavy field $\chi$; the problem is that this produces a sextic term with the wrong sign, resulting in a potential that is unbounded below and allowing for infinitely negative energy density. To stabilize this, one would need to add additional non-renormalizable terms, thus creating more problems. The use of $w=1/2$ is thus problematic and we avoid it. Fortunately, we find that $w=1/2$ is seemingly not needed to mimic the $H(z)$ favored by DESI.
@@ VII Appendix
VII.1 Matrix Element for Tree-Level Scattering Cross Section
Here we include a closed-form expression for the matrix element used in plotting the SIDM cross section. This was produced from the $t$ and $u$-channel Feynman diagrams via a manual computation. For compactness, we define $C_{1}$, $C_{2}$, and $C_{3}$ as the following quantities. We use p for the momentum, $\mathrm{E}_{\mathrm{cm}}$ for the energy in the center-of-mass frame, $\theta$ for the scattering angle, $m_{h}$ for the mass of the scalar mediator, and $m_{\alpha}$ for the mass of the vector field.
@@ VII.2 Tree-Level Decay Rate
After symmetry breaking, we get interaction terms of the form $\frac{g_{\rm{AB}}\nu}{\sqrt{2}}\rm{A}\rm{B}^{2}$, where $\nu$ is the VEV (vacuum expectation value) of the dark Higgs. We use $\nu\approx 2.725(1+z_{\rm{BREAK}})k_{B}$, where $k_{B}$ is the Boltzmann constant, $z_{\rm{BREAK}}$ is the redshift at symmetry breaking, and the CMB temperature at $z=0$ is 2.725 Kelvin. The decay rate for A $\to$ 2B can be computed from the particle masses, $m_{\rm{A}}$ and $m_{\rm{B}}$, and the coupling constant, $g_{\rm{AB}}$, as shown below.
@@
Here we have $c=1$, $\nu$ in units of eV, masses in units of eV, $\hbar\approx 6.58\times 10^{-16}$ eV $\cdot$ sec, and $\Gamma$ in units of $\rm{sec}^{-1}$. We see that our desired decay rates can be achieved using the following parameter values. These should be regarded as example values only, as there is a fair bit of flexibility. One could increase these coupling constants, resulting in much larger decay rates, but the effect on H(z) would be similar. The main motivation for not working with decay rates greater than $10^{-15}\rm{sec}^{-1}$ is that this would necessitate a smaller time step in the numerical integration of the fluid equations.
Note that we have chosen masses such that the daughter particle masses are slightly less than half the masses of the parent particles; this ensures the decay products are cold. We also note that since the dark Higgs is very light, and the mass of B is very close to half that of A, we can neglect any energy transfer from A to $h$ via $\frac{g_{\rm{AB}}h}{\sqrt{2}}\rm{A}\rm{B}^{2}$, where $h$ is the radial mode of the dark Higgs after symmetry breaking.
Field Mass
$\rm{A}$ 100 keV
$\rm{B}$ 49.999 keV
$\rm{C}$ 24.999 keV
$\rm{D}$ 12.499 keV
Table 2: We provide example values for dark matter masses that yield $\Gamma=10^{-15}\rm{sec}^{-1}$. We note that the masses of the daughter particles are chosen to be slightly less than half those of the parent particles so that the decay products are cold and the energy transfer from the parent to the radial mode of the dark Higgs is minimal.
Coupling Constant Value
$g_{\rm AB}$ $2.4\times 10^{-8}$
$g_{\rm BC}$ $2.3\times 10^{-8}$
$g_{\rm CD}$ $2.9\times 10^{-8}$
Table 3: We provide example values for decay coupling constants that yield $\Gamma=10^{-15}~\mathrm{sec}^{-1}$.

### v2

@@ IV.2 SIDM Cross Section
We include the closed-form expression for $|\rm{M}|^{2}$ in the Appendix. We employ the tree-level cross section instead of partial wave analysis or the simpler Born approximation because the tree-level cross section captures physics that the other two do not. While partial wave analysis has been used to capture non-perturbative effects for a Yukawa interaction [63], we have vectors interacting via a scalar, rather than fermions. Massive vector fields differ from fermions because unlike fermions, they are bosons with three distinct polarizations, which affects the cross section. The Born approximation offers a simpler method for computing a cross section from a Yukawa potential, though it neglects the $u$-channel diagram, as well as the interference between $t$ and $u$-channels [64], which our calculation incorporates. (See Figure 1 for a depiction of the $t$ and $u$-channel processes.) One can view neglecting the $u$-channel as a simplifying approximation, though it dramatically changes $d\sigma/d\Omega$. A process that includes only the $t$-channel yields a differential cross section that peaks at $\theta=0$ but not $\theta=\pi$, whereas the true differential cross section peaks at both $\theta=0$ and $\theta=\pi$, producing a U-shaped plot. This can in turn yield a noticeable impact on the momentum transfer cross section [64], so it is preferable to use the full tree-level cross section, when possible, even though this is more complicated to compute.
@@ V Conclusion
We find it to be very difficult to produce a natural, fully renormalizable model that can realize the behavior desired by Chen and Loeb, though we are able to produce a more natural effective field theory that can generate similar behavior. This model also naturally yields a possible SIDM candidate as a consequence of gauge invariance. In a following work, we plan to replace the cosmological constant with a quintessence field not coupled to the dark matter. Future work could also fit the model parameters to DESI data, study effects on large-scale structure, consider production mechanisms for the dark matter, and more thoroughly investigate the SIDM candidate. It would be interesting to incorporate the tree-level differential cross section into N-body codes to correctly incorporate scattering that is both velocity-dependent, and anisotropic. While there are a few possibilities for production, one mechanism could involve inflaton decay [65], or coupling a dark vector field to the inflaton [66].
We emphasize that even if future work finds our fiducial model to be disfavored, this study is still valuable for studying the space of effective field theories that have the potential to realize the behavior desired by Chen and Loeb, which is relevant for finding a pathway towards explaining the DESI trends.
@@ VII Appendix
VII.1 S-Matrix Amplitude for Tree-Level Scattering Cross Section
Here we include a closed-form expression for the S-matrix amplitude used in plotting the SIDM cross section. This was produced from the $t$ and $u$-channel Feynman diagrams via a manual computation. For compactness, we define $C_{1}$, $C_{2}$, and $C_{3}$ as the following quantities. We use p for the momentum, $\mathrm{E}_{\mathrm{cm}}$ for the energy in the center-of-mass frame, $\theta$ for the scattering angle, $m_{h}$ for the mass of the scalar mediator, and $m_{\alpha}$ for the mass of the vector field.
@@ VII.2 Tree-Level Decay Rate
After symmetry breaking, we get interaction terms of the form $\frac{g_{\rm{AB}}\nu}{\sqrt{2}}\rm{A}\mathrm{B}^{\dagger}\mathrm{B}$, where $\nu$ is the VEV (vacuum expectation value) of the dark Higgs. The decay rate for A $\to$ $\mathrm{B}^{\dagger}\mathrm{B}$ can be computed from the particle masses, $m_{\rm{A}}$ and $m_{\rm{B}}$, and the coupling constant, $g_{\rm{AB}}$, as shown below, with $c=1$, $\nu$ in units of eV, masses in units of eV, $\hbar\approx 6.58\times 10^{-16}$ eV $\cdot$ sec, and $\Gamma$ in units of $\rm{sec}^{-1}$. The mass of A (and by extension, the masses of B, C, and D) depends in part on the production mechanism for the dark matter. Future work can determine masses through a fit to observations after consideration of production mechanisms.
@@
In our analysis, we neglect any energy transfer from A to $h$ via $\frac{g_{\rm{A}\rm{B}}h}{\sqrt{2}}\rm{A}\mathrm{B}^{\dagger}\mathrm{B}$, where $h$ is the massive radial mode of the dark Higgs after symmetry breaking. A more detailed analysis could consider this for the decay of A, as well as similarly for decays of B and C.